Find the degree measure of one interior and one exterior.

( 20 sides polygon)

Answer by stanbon(75887)

Find the degree measure of one interior and one exterior.
( 20 sides polygon)
--------------------------
The sum of the exterior angles is 360 degrees.
---
Each exterior angle has measure = 360/20 = 18 degrees
----
Each interior angle is the supplement of an exterior angle.
----
Each interior angle is 180-18 = 162 degrees
===============================================
Cheers,
Stan H.